# Advances in the Shannon Capacity of Graphs

### v1

@@ 2.3 The Lovász function of graphs
Definition 2.26 (Lovász function).
Let $\mathsf{G}$ be a simple graph of order $n$. The Lovász function of $\mathsf{G}$ is defined as
@@
An orthonormal representation of the pentagon $\operatorname{\mathsf{C}}_{5}$, along with its handle $\mathbf{c}$, is shown in Figure 1.
@@
The Lovász $\vartheta$-function can be expressed as a solution of an SDP problem. To that end, let ${\bf{A}}=(A_{i,j})$ be the $n\times n$ adjacency matrix of $\mathsf{G}$ with $n\triangleq|\operatorname{\mathsf{V}}(\mathsf{G})|$. The Lovász $\vartheta$-function $\vartheta(\mathsf{G})$ can be expressed by the following convex optimization problem:
@@
In light of (42), one can ask if the Schrijver $\vartheta$-function can serve as a better upper bound for the Shannon capacity. However, in a recent paper by the second author [30], it was shown that the Schrijver $\vartheta$-function need note be an upper bound on the Shannon capacity, by providing a graph whose Shannon capacity is strictly larger than the Schrijver $\vartheta$-function. Lastly, the Schrijver $\vartheta$-function is an upper bound on the independence number, so it can be used for constructing better bounds on the independence number.
@@
[11] Let $(\mathbf{v}_{1},\cdots,\mathbf{v}_{n})$ range over all orthonormal representations of $\overline{\mathsf{G}}$ and $\mathbf{d}$ over all unit vectors. Then
@@
Next, we provide several properties of the Lovász number, regarding graph operations and subgraphs.
@@
The following result was first stated and proved by Knuth (Section 18 of [16]). We suggest an alternative elementary proof in Appendix A.
@@ 2.4 Concluding preliminaries
We next provide a useful lower bound for the fractional independence number and the fractional chromatic number (see [32, Proposition 3.1.1]).
@@
$\alpha_{\mathrm{f}}(\mathsf{G})$ $\geq\frac{n}{\operatorname{\omega}(\mathsf{G})},$
$\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$ $\geq\frac{n}{\alpha(\mathsf{G})}.$
@@
$\alpha(\mathsf{G})$ $=\binom{n-1}{r-1}$,
$\operatorname{\omega}(\mathsf{G})$ $=\bigg\lfloor\frac{n}{r}\bigg\rfloor,$
$\Theta(\mathsf{G})$ $=\binom{n-1}{r-1}$,
$\vartheta(\mathsf{G})$ $=\binom{n-1}{r-1}$,
$\vartheta(\overline{\mathsf{G}})$ $=\frac{n}{r},$
$\alpha_{\mathrm{f}}(\mathsf{G})$ $=\frac{\binom{n}{r}}{\lfloor\frac{n}{r}\rfloor},$
$\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$ $=\frac{n}{r},$
$\operatorname{\chi}(\mathsf{G})$ $=n-2r+2,$
$\operatorname{\sigma}(\mathsf{G})$ $=\Bigg\lceil\frac{\binom{n}{r}}{\big\lfloor\frac{n}{r}\big\rfloor}\Bigg\rceil.$
@@
Let $\ell\geq 1$. Then the path graph $\operatorname{\mathsf{P}}_{\ell}$ is a universal graph with parameters
@@
Let $k\geq 4$. Then the cycle graph $\operatorname{\mathsf{C}}_{k}$ has
@@
    $\alpha(\operatorname{\mathsf{C}}_{k})=\Theta(\operatorname{\mathsf{C}}_{k})=\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{2}.$
@@
    $\alpha(\operatorname{\mathsf{C}}_{k})=\left\lfloor\frac{k}{2}\right\rfloor\;\;\;;\;\;\;\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{1+\sec\frac{\pi}{k}}.$
@@
If $\mathsf{G}$ is a universal graph, then by Theorems 2.6 and 2.21,
$\Theta(\mathsf{G})=\alpha(\mathsf{G}).$
Remark $2.$
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}(n,r)$ be a Kneser graph with $n\geq 2r$, then $\mathsf{G}$ is universal if and only if $r|n$.
@@
$\alpha_{\mathrm{f}}(\mathsf{G})=\frac{\binom{n}{r}}{\lfloor\frac{n}{r}\rfloor}.$
For $\mathsf{G}$ to be a universal graph, the equality $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ has to hold (see Theorem 2.6), thus, $\mathsf{G}$ is universal if and only if:
$\binom{n-1}{r-1}=\frac{\binom{n}{r}}{\lfloor\frac{n}{r}\rfloor},$
which holds if and only if $r|n$, as required. ∎
@@
The Petersen graph $\mathsf{G}=\operatorname{\mathsf{K}}(5,2)$ has $\alpha(\mathsf{G})=\Theta(\mathsf{G})$ but is not universal.
@@ 3 The Shannon capacity of polynomials in graphs
Shannon conjectured that for any two graphs $\mathsf{G}$ and $\mathsf{H}$, the capacity of their disjoint union is equal to the sum of their individual capacities [1]. This conjecture was later disproved by Alon [7], who showed that if $\mathsf{G}$ is the Schläfli graph, then
$\Theta(\mathsf{G})+\Theta(\overline{\mathsf{G}})\leq 10<2\sqrt{27}\leq\Theta(\mathsf{G}+\overline{\mathsf{G}}).$
In this section, we provide sufficient conditions on a sequence of graphs, such that the capacity of their disjoint union is equal to the sum of their individual capacities. The results of this section rely on a recent result by Schrijver [24]. Due to its importance, it is provided with a proof in the preliminaries (see Theorem 2.12), stating that for any two simple graphs $\mathsf{G}$ and $\mathsf{H}$,
$\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})\iff\Theta(\mathsf{G}+\mathsf{H})=\Theta(\mathsf{G})+\Theta(\mathsf{H}).$
This result was independently proved by Wigderson and Zuiddam [22], with credit to Holzman for personal communications.
Define the set $\mathbb{N}[x_{1},\ldots,x_{\ell}]$ to be the set of all nonzero polynomials of variables $x_{1},\ldots,x_{\ell}$ with nonnegative integral coefficients. A polynomial in graphs, $p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})$, is defined such that the plus operation of graphs stands for their disjoint union and a product of graphs stands for their strong product.
@@
By a corollary of the duality theorem (Theorem 2.12), which was proven by Schrijver [24, Theorem 3], the following surprising result holds.
@@
By Lemma 3.1, we focus on finding sufficient conditions for the satisfiability of (52). In this section, we present two main results that provide such sufficient conditions, discuss the differences between them, and illustrate their application. We start by providing the following result.
@@
By Theorems 2.10 and 2.11,
$p(\Theta(\mathsf{G}_{1}),\ldots,\Theta(\mathsf{G}_{\ell}))\leq\Theta(p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})),$
and by Theorems 2.15 and 2.16,
@@
$=p(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell})).$
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, with $\ell\in\mathbb{N}$. If $\Theta(\mathsf{G}_{i})=\vartheta(\mathsf{G}_{i})$ for every $i\in[\ell]$, then
$\Theta(p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}))=p(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell}))$ (54)
for every $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$.
@@
$p(\Theta(\mathsf{G}_{1}),\ldots,\Theta(\mathsf{G}_{\ell}))=p(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell})).$
Thus, by Theorem 3.1, equality (54) holds. ∎
We next provide our first sufficient condition.
@@
By the assumption and Theorem 2.16,
$\vartheta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})$ $=\vartheta(\mathsf{G}_{1})+\ldots+\vartheta(\mathsf{G}_{\ell})$
$=\Theta(\mathsf{G}_{1})+\ldots+\Theta(\mathsf{G}_{\ell})$
$\leq\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})$
$\leq\vartheta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell}),$
which gives
$\Theta(\mathsf{G}_{1}+\mathsf{G}_{2}+\ldots+\mathsf{G}_{\ell})=\Theta(\mathsf{G}_{1})+\Theta(\mathsf{G}_{2})+\ldots+\Theta(\mathsf{G}_{\ell}).$
@@
If $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ are all Kneser graphs or self-complementary vertex-transitive graphs, then by Theorem 3.2, equality (51) holds for every polynomial $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$. In particular, if $\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i})$ (with $n_{i}\geq 2r_{i}$) for every $i\in[\ell]$, then
$\Theta(\operatorname{\mathsf{K}}(n_{1},r_{1})+\ldots+\operatorname{\mathsf{K}}(n_{\ell},r_{\ell}))=\binom{n_{1}-1}{r_{1}-1}+\ldots+\binom{n_{\ell}-1}{r_{\ell}-1}$.
Similarly, if $\mathsf{G}_{i}$ is a self-complementary and vertex-transitive graph for every $i\in[\ell]$, then by [34, Theorem 3.26]
$\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})=\sqrt{n_{1}}+\ldots+\sqrt{n_{\ell}},$
@@
$\Theta(\mathsf{G}_{1}\boxtimes\mathsf{G}_{2})$ $=\lim_{k\to\infty}\sqrt[k]{\alpha((\mathsf{G}_{1}\boxtimes\mathsf{G}_{2})^{k})}$ (57)
$\leq\lim_{k\to\infty}\sqrt[k]{\alpha_{\mathrm{f}}(\mathsf{G}_{1}^{k})\alpha(\mathsf{G}_{2}^{k})}$ (58)
@@
Combining inequalities (56) and (57) gives
@@
$\leq\alpha_{\mathrm{f}}(\mathsf{G}_{1})\,\Theta(\mathsf{G}_{2}).$
@@
$\Theta(\mathsf{G}_{1})\,\Theta(\mathsf{G}_{2})=\Theta(\mathsf{G}_{1}\boxtimes\mathsf{G}_{2}),$
@@
$\Theta\left(\sum_{i=1}^{\ell}{\mathsf{G}_{i}}\right)=\Theta(\mathsf{G}_{1})+\Theta\left(\sum_{i=2}^{\ell}{\mathsf{G}_{i}}\right)=\ldots=\sum_{i=1}^{\ell}{\Theta(\mathsf{G}_{i})}.$
@@
It is worth noting that our sufficient conditions for equality (51) to hold are less restrictive than those hinted in Shannon’s paper [1], which are next presented. The following theorem is derived directly from Lemma 3.2 and the sandwich theorem. We provide the original proof in Appendix B.
@@
Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs with $\alpha(\mathsf{G}_{1})=\operatorname{\chi}(\overline{\mathsf{G}}_{1})$. Then
$\Theta(\mathsf{G}_{1}+\mathsf{G}_{2})=\Theta(\mathsf{G}_{1})+\Theta(\mathsf{G}_{2}).$
@@
$\alpha(\mathsf{G})\leq\Theta(\mathsf{G})\leq\alpha_{\mathrm{f}}(\mathsf{G})\leq\operatorname{\chi}(\overline{\mathsf{G}}).$
@@
Let $\mathsf{G}=\overline{\operatorname{\mathsf{K}}(n,r)}$ be the complement of a Kneser graph with $n>2r>1$ and $r|n$. Claim: the graph $\mathsf{G}$ has
$\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})<\operatorname{\chi}(\overline{\mathsf{G}}).$
The invariants of the Kneser graph and its complement are known and are presented in Theorem 2.22, according to that theorem,
$\alpha(\mathsf{G})$ $=\operatorname{\omega}(\operatorname{\mathsf{K}}(n,r))=\frac{n}{r}$
$\alpha_{\mathrm{f}}(\mathsf{G})$ $=\frac{n}{r}$
$\operatorname{\chi}(\overline{\mathsf{G}})$ $=\operatorname{\chi}(\operatorname{\mathsf{K}}(n,r))=n-2r+2,$
since the independence number and the fractional independence number are equal to $\frac{n}{r}$ then $\Theta(\mathsf{G})=\frac{n}{r}$. Hence, because $n>2r$, we have
$\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})=\frac{n}{r}<n-2r+2=\operatorname{\chi}(\overline{\mathsf{G}}),$
as required.
@@
$\Theta(\mathsf{G})\leq\vartheta(\mathsf{G})\leq\alpha_{\mathrm{f}}(\mathsf{G}),$
so the condition $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ is less restrictive than the condition $\Theta(\mathsf{G})=\vartheta(\mathsf{G})$. Consequently, Theorem 3.3 imposes a stronger condition than the one in Theorem 3.2 on $\ell-1$ of the graphs, while no condition is imposed on the $\ell$-th graph. The following examples clarify the issue.
Example 3.3.
@@
$\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i})\,\,\,\,\,(n_{i}\geq 2r_{i}).$
For these graphs we have
$\alpha(\mathsf{G}_{i})=\Theta(\mathsf{G}_{i})=\vartheta(\mathsf{G}_{i})=\binom{n_{i}-1}{r_{i}-1}\;\;\;,\;\;\;\forall{i}\in[\ell].$
On the other hand, by Theorem 2.22,
$\alpha_{\mathrm{f}}(\mathsf{G}_{i})=\frac{\binom{n_{i}}{r_{i}}}{\lfloor\frac{n_{i}}{r_{i}}\rfloor}.$ (60)
If $r_{i}\nmid n_{i}$, then it follows that
@@
Let the parameters $n_{1},\ldots,n_{\ell}\,\,,\,\,r_{1},\ldots,r_{\ell}$ be selected such that $r_{i}\nmid n_{i}$ for at least two of the graphs $\{\mathsf{G}_{i}\}_{i=1}^{\ell}$, we have $\alpha_{\mathrm{f}}(\mathsf{G}_{i})>\Theta(\mathsf{G}_{i})$, which then violates the satisfiability of the sufficient conditions in Theorem 3.3. Hence, in that case
$\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})=\sum_{i=1}^{\ell}\binom{n_{i}-1}{r_{i}-1}$
by Theorem 3.2, whereas this equality is not implied by Theorem 3.3.

### v2

@@ 2.3 The Lovász $\vartheta$-function of graphs
Definition 2.26 (Lovász $\vartheta$-function).
Let $\mathsf{G}$ be a simple graph of order $n$. The Lovász $\vartheta$-function of $\mathsf{G}$ is defined as
@@
An orthonormal representation of the pentagon $\operatorname{\mathsf{C}}_{5}$, along with its handle $\mathbf{c}$, is shown in Figure 1. This figure is reproduced, with permission of the author [34], by combining Figures 1.3 and 1.4 in [12].
@@
The Lovász $\vartheta$-function can be expressed as a solution of a semidefinite programming (SDP) problem. To that end, let ${\bf{A}}=(A_{i,j})$ be the $n\times n$ adjacency matrix of $\mathsf{G}$ with $n\triangleq|\operatorname{\mathsf{V}}(\mathsf{G})|$. The Lovász $\vartheta$-function $\vartheta(\mathsf{G})$ can be expressed by the following convex optimization problem:
@@
In light of (41), one may ask whether the Schrijver $\vartheta$-function can serve as a tighter upper bound on the Shannon capacity of graphs. However, a recent paper by the second author [35] shows that the Schrijver $\vartheta$-function does not, in general, provide an upper bound on the Shannon capacity: an explicit graph is constructed whose Shannon capacity is strictly larger than its Schrijver $\vartheta$-value. Nevertheless, the Schrijver $\vartheta$-function is an upper bound on the independence number, and can therefore be used to derive an improved bound on the independence number.
@@
[11] Let $(\mathbf{v}_{1},\cdots,\mathbf{v}_{n})$ range over all orthonormal representations of $\overline{\mathsf{G}}$ and $\mathbf{d}$ over all unit vectors. Then,
@@
Next, we provide several properties of the Lovász $\vartheta$-function, regarding graph operations and subgraphs.
@@
The following result was first stated and proved by Knuth (Section 18 of [17]). We suggest an alternative elementary proof in Appendix B.
@@
The following result is an easy consequence of Definition 2.26, and it is an analog of Theorem 2.14.
@@ 2.4 Concluding preliminaries
We next provide a useful lower bound for the fractional independence number and the fractional chromatic number (see [37, Proposition 3.1.1]).
@@
$\alpha_{\mathrm{f}}(\mathsf{G})$ $\geq\frac{n}{\operatorname{\omega}(\mathsf{G})},$ (48)
$\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$ $\geq\frac{n}{\alpha(\mathsf{G})}.$ (49)
@@
$\alpha(\mathsf{G})$ $=\binom{n-1}{r-1}$, (51)
$\operatorname{\omega}(\mathsf{G})$ $=\bigg\lfloor\frac{n}{r}\bigg\rfloor,$ (52)
$\Theta(\mathsf{G})$ $=\binom{n-1}{r-1}$, (53)
$\vartheta(\mathsf{G})$ $=\binom{n-1}{r-1}$, (54)
$\vartheta(\overline{\mathsf{G}})$ $=\frac{n}{r},$ (55)
$\alpha_{\mathrm{f}}(\mathsf{G})$ $=\frac{\binom{n}{r}}{\bigl\lfloor\frac{n}{r}\bigr\rfloor},$ (56)
$\operatorname{\chi}_{\mathrm{f}}(\mathsf{G})$ $=\frac{n}{r},$ (57)
$\operatorname{\chi}(\mathsf{G})$ $=n-2r+2,$ (58)
$\operatorname{\sigma}(\mathsf{G})$ $=\Bigg\lceil\frac{\binom{n}{r}}{\big\lfloor\frac{n}{r}\big\rfloor}\Bigg\rceil.$ (59)
@@
The path $\operatorname{\mathsf{P}}_{\ell}$, for $\ell\in\mathbb{N}$, is a universal graph with parameters
@@
Let $k\geq 4$. Then, the cycle graph $\operatorname{\mathsf{C}}_{k}$ has
@@
    $\alpha(\operatorname{\mathsf{C}}_{k})=\Theta(\operatorname{\mathsf{C}}_{k})=\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{2}.$ (61)
@@
    $\alpha(\operatorname{\mathsf{C}}_{k})=\left\lfloor\frac{k}{2}\right\rfloor,\quad\vartheta(\operatorname{\mathsf{C}}_{k})=\frac{k}{1+\sec\frac{\pi}{k}}.$ (62)
@@
If $\mathsf{G}$ is a universal graph, then by Theorems 2.6 and 2.22,
$\Theta(\mathsf{G})=\alpha(\mathsf{G}).$ (63)
Remark 2.1.
@@
Let $\mathsf{G}=\operatorname{\mathsf{K}}(n,r)$ be a Kneser graph with $n\geq 2r$. Then, $\mathsf{G}$ is universal if and only if $r\mid n$.
@@
$\alpha_{\mathrm{f}}(\mathsf{G})=\frac{\binom{n}{r}}{\lfloor\frac{n}{r}\rfloor}.$ (64)
For $\mathsf{G}$ to be a universal graph, the equality $\alpha(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ has to hold (see Theorem 2.6), thus, $\mathsf{G}$ is universal if and only if
$\binom{n-1}{r-1}=\frac{\binom{n}{r}}{\lfloor\frac{n}{r}\rfloor},$ (65)
which holds if and only if $r\mid n$, as required. ∎
@@
The Petersen graph $\mathsf{G}=\operatorname{\mathsf{K}}(5,2)$ satisfies $\alpha(\mathsf{G})=\Theta(\mathsf{G})$, yet it is not universal by Corollary 2.1.
@@ 3 The Shannon capacity of polynomials in graphs
The Shannon capacity of the disjoint union of graphs is at least the sum of their individual capacities [1]; however, it may in fact be significantly larger [7]. This phenomenon is strikingly demonstrated by a result of Alon [7, Theorem 1.1], which we restate here as Theorem 2.11. Moreover, Alon showed in [7, Theorem 1.2] that if we let $\mathsf{G}$ be the Schläfli graph, then
$\Theta(\mathsf{G})+\Theta(\overline{\mathsf{G}})\leq 10<2\sqrt{27}\leq\Theta(\mathsf{G}+\overline{\mathsf{G}}).$ (66)
We note that the simple counterexample constructed earlier by Haemers [15] plays a key role in this argument and was subsequently employed by Alon [7] in establishing (66).
We provide sufficient conditions on a sequence of graphs under which the Shannon capacity of their disjoint union equals the sum of their individual capacities. Our results rely on a recent result by Schrijver [28], which we include with proof due to its central role (see Theorem 2.12). It states that for any two simple graphs $\mathsf{G}$ and $\mathsf{H}$,
$\Theta(\mathsf{G}\boxtimes\mathsf{H})=\Theta(\mathsf{G})\,\Theta(\mathsf{H})\iff\Theta(\mathsf{G}+\mathsf{H})=\Theta(\mathsf{G})+\Theta(\mathsf{H}).$ (67)
This result was independently proved by Wigderson and Zuiddam [23], with credit to Holzman for personal communication.
Let $\mathbb{N}[x_{1},\ldots,x_{\ell}]$ be the set of nonzero polynomials in the variables $x_{1},\ldots,x_{\ell}$ with nonnegative integer coefficients. A polynomial in graphs, $p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})$, is defined by interpreting graph addition as disjoint union and graph multiplication as strong product.
@@
By a corollary of the duality theorem (Theorem 2.12), which was proved by Schrijver [28], the following surprising result holds.
@@
By Lemma 3.1 (see [28, Theorem 3]), we focus on finding sufficient conditions for the satisfiability of (69). In this section, we present two main results that provide such sufficient conditions, discuss the differences between them, and illustrate their application. We start by providing the following result.
@@
By Theorems 2.9 and 2.10,
$p(\Theta(\mathsf{G}_{1}),\ldots,\Theta(\mathsf{G}_{\ell}))\leq\Theta(p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell})),$ (71)
and by Theorems 2.16 and 2.17,
@@
$=p(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell})).$ (72)
@@
Let $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ be simple graphs, and let $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$ with $\ell\in\mathbb{N}$. If $\Theta(\mathsf{G}_{i})=\vartheta(\mathsf{G}_{i})$ for every $i\in[\ell]$, then
$\Theta(p(\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}))=p\bigl(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell})\bigr).$ (73)
@@
$p(\Theta(\mathsf{G}_{1}),\ldots,\Theta(\mathsf{G}_{\ell}))=p\bigl(\vartheta(\mathsf{G}_{1}),\ldots,\vartheta(\mathsf{G}_{\ell})\bigr).$ (74)
Thus, equality (73) holds by Theorem 3.1. ∎
The next result readily follows.
@@
This holds by Corollary 3.1 upon replacing $\Theta(\mathsf{G}_{i})$ with $\vartheta(\mathsf{G}_{i})$ for all $i\in[\ell]$. ∎
@@
If $\mathsf{G}_{1},\ldots,\mathsf{G}_{\ell}$ are all Kneser graphs or self-complementary vertex-transitive graphs, then by Theorem 3.2, equality (68) holds for every polynomial $p\in\mathbb{N}[x_{1},\ldots,x_{\ell}]$. In particular, if $\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i})$ (with $n_{i}\geq 2r_{i}$) for every $i\in[\ell]$, then by Theorem 2.23
$\Theta(\operatorname{\mathsf{K}}(n_{1},r_{1})+\ldots+\operatorname{\mathsf{K}}(n_{\ell},r_{\ell}))=\binom{n_{1}-1}{r_{1}-1}+\ldots+\binom{n_{\ell}-1}{r_{\ell}-1}$. (75)
Similarly, if $\mathsf{G}_{i}$ is self-complementary and vertex-transitive for all $i\in[\ell]$, then by [39, Theorem 3.26]
$\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})=\sqrt{n_{1}}+\ldots+\sqrt{n_{\ell}},$ (76)
@@
$\Theta(\mathsf{G}_{1}\boxtimes\mathsf{G}_{2})$ $=\lim_{k\to\infty}\sqrt[k]{\alpha((\mathsf{G}_{1}\boxtimes\mathsf{G}_{2})^{k})}$
$\leq\lim_{k\to\infty}\sqrt[k]{\alpha_{\mathrm{f}}(\mathsf{G}_{1}^{k})\,\alpha(\mathsf{G}_{2}^{k})}$
@@
where (79) holds since the equality $\alpha_{\mathrm{f}}(\mathsf{G}_{1})=\sqrt[k]{\alpha_{\mathrm{f}}(\mathsf{G}_{1}^{k})}$ is satisfied for all $k\in\mathbb{N}$, and by definition $\Theta(\mathsf{G}_{2})=\underset{k\to\infty}{\lim}\sqrt[k]{\alpha(\mathsf{G}_{2}^{k})}$. Combining inequalities (78) and (79) gives
@@
$\leq\alpha_{\mathrm{f}}(\mathsf{G}_{1})\,\Theta(\mathsf{G}_{2}).$ (80)
@@
$\Theta(\mathsf{G}_{1})\,\Theta(\mathsf{G}_{2})=\Theta(\mathsf{G}_{1}\boxtimes\mathsf{G}_{2}),$ (81)
@@
$\Theta\left(\sum_{i=1}^{\ell}{\mathsf{G}_{i}}\right)=\Theta(\mathsf{G}_{1})+\Theta\left(\sum_{i=2}^{\ell}{\mathsf{G}_{i}}\right)=\ldots=\sum_{i=1}^{\ell}{\Theta(\mathsf{G}_{i})}.$ (82)
@@
It is worth noting that our sufficient conditions for equality (68) to hold are less restrictive than those hinted in Shannon’s paper [1], which are next presented. The following theorem is derived directly from Lemma 3.2 and the sandwich theorem. We provide the original proof in Appendix C.
@@
Let $\mathsf{G}_{1}$ and $\mathsf{G}_{2}$ be simple graphs with $\alpha(\mathsf{G}_{1})=\operatorname{\chi}(\overline{\mathsf{G}}_{1})$. Then,
@@
$\alpha(\mathsf{G})\leq\Theta(\mathsf{G})\leq\alpha_{\mathrm{f}}(\mathsf{G})\leq\operatorname{\chi}(\overline{\mathsf{G}}).$ (84)
@@
Let $\mathsf{G}=\overline{\operatorname{\mathsf{K}}(n,r)}$ be the complement of a Kneser graph where $n>2r$, $r>1$, and $r\mid n$. It is claimed that the graph $\mathsf{G}$ satisfies
$\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})<\operatorname{\chi}(\overline{\mathsf{G}}).$ (85)
Indeed, the invariants of the Kneser graph and its complement are known, and by Theorem 2.23 (since $\alpha(\mathsf{G})=\operatorname{\omega}(\overline{\mathsf{G}})$, where $\overline{\mathsf{G}}=\operatorname{\mathsf{K}}(n,r)$ with $r$ and $n$ as above)
$\alpha(\mathsf{G})$ $=\frac{n}{r},$ (86)
$\alpha_{\mathrm{f}}(\mathsf{G})$ $=\frac{n}{r},$ (87)
$\operatorname{\chi}(\overline{\mathsf{G}})$ $=n-2r+2.$ (88)
Since the independence number and the fractional independence number coincide and are both equal to $\frac{n}{r}$ then $\Theta(\mathsf{G})=\frac{n}{r}$. Hence, the assumption $n>2r$ implies that
$\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})=\frac{n}{r}<n-2r+2=\operatorname{\chi}(\overline{\mathsf{G}}).$ (89)
@@
$\Theta(\mathsf{G})\leq\vartheta(\mathsf{G})\leq\alpha_{\mathrm{f}}(\mathsf{G}),$ (90)
so the condition $\Theta(\mathsf{G})=\alpha_{\mathrm{f}}(\mathsf{G})$ implies that the equality $\Theta(\mathsf{G})=\vartheta(\mathsf{G})$ holds. Hence, Theorem 3.3 imposes a stronger condition than the one in Theorem 3.2 on $\ell-1$ of the graphs, while no condition is imposed on the $\ell$-th graph.
The next two examples (Examples 3.3 and 3.4) show that neither Theorem 3.2 nor Theorem 3.3 implies the other.
Example $3.$
@@
$\mathsf{G}_{i}=\operatorname{\mathsf{K}}(n_{i},r_{i}),\quad n_{i}\geq 2r_{i}.$ (91)
For these non-empty graphs, by Theorem 2.23,
$\alpha(\mathsf{G}_{i})=\Theta(\mathsf{G}_{i})=\vartheta(\mathsf{G}_{i})=\binom{n_{i}-1}{r_{i}-1},\quad\forall\,i\in[\ell],$ (92)
and
$\alpha_{\mathrm{f}}(\mathsf{G}_{i})=\frac{\binom{n_{i}}{r_{i}}}{\lfloor\frac{n_{i}}{r_{i}}\rfloor},\quad\forall\,i\in[\ell].$ (93)
If $r_{i}\nmid n_{i}$, then it follows from (92) and (93) that
@@
Let the parameters $n_{1},\ldots,n_{\ell}$ and $r_{1},\ldots,r_{\ell}$ be chosen such that $r_{i}\nmid n_{i}$ for at least two of the Kneser graphs $\{\mathsf{G}_{i}\}_{i=1}^{\ell}$. For each such index $i\in[\ell]$, we have $\alpha_{\mathrm{f}}(\mathsf{G}_{i})>\Theta(\mathsf{G}_{i})$, which violates the sufficient conditions of Theorem 3.3. By (92), it follows from Theorem 3.2 that
$\Theta(\mathsf{G}_{1}+\ldots+\mathsf{G}_{\ell})=\sum_{i=1}^{\ell}\binom{n_{i}-1}{r_{i}-1}$, (95)
whereas equality (95) does not follow from Theorem 3.3.